# Curriculum Negative Mining For Temporal Networks

### v1

@@ Abstract
Temporal networks are effective in capturing the evolving interactions of networks over time, such as social networks and e-commerce networks. In recent years, researchers have primarily concentrated on developing specific model architectures for Temporal Graph Neural Networks (TGNNs) in order to improve the representation quality of temporal nodes and edges. However, limited attention has been given to the quality of negative samples during the training of TGNNs. When compared with static networks, temporal networks present two specific challenges for negative sampling: positive sparsity and positive shift. Positive sparsity refers to the presence of a single positive sample amidst numerous negative samples at each timestamp, while positive shift relates to the variations in positive samples across different timestamps. To robustly address these challenges in training TGNNs, we introduce Curriculum Negative Mining (CurNM), a model-aware curriculum learning framework that adaptively adjusts the difficulty of negative samples. Within this framework, we first establish a dynamically updated negative pool that balances random, historical, and hard negatives to address the challenges posed by positive sparsity. Secondly, we implement a temporal-aware negative selection module that focuses on learning from the disentangled factors of recently active edges, thus accurately capturing shifting preferences. Extensive experiments on 12 datasets and 3 TGNNs demonstrate that our method outperforms baseline methods by a significant margin. Additionally, thorough ablation studies and parameter sensitivity experiments verify the usefulness and robustness of our approach. Our code is available at https://github.com/zziyue83/CurNM.
@@
[2]organization=Big Graph Center, School of Computer and Computing Science, Hangzhou City University,city=Hangzhou, state=Zhejiang, country=China
@@ 1 Introduction
Similar to static graph models, TGNNs [24, 33] employ negative sampling techniques to prevent the models from being overwhelmed by a disproportionally high number of negative over positive samples and enhance the models’ efficiency. The most straightforward approaches often adopted are random sampling and recent sampling [24, 33]. However, temporal networks introduce two unique challenges that particularly undermine these approaches: positive sparsity and positive shift. Positive sparsity emerges as the inclusion of a temporal dimension significantly increases the negative-to-positive ratio by the number of timestamps, compared to static graphs. Consequently, random sampling in this context tends to produce a multitude of irrelevant and easily distinguishable negative samples, offering little to no value. Positive shift, on the other hand, occurs as the relationships between nodes change over time; for instance, in recommendation systems, users’ preferences are likely to evolve [16]. Ignoring this factor may result in training models on outdated information, particularly in datasets spanning long periods.
@@
To address these challenges, we introduce the CurNM method. Overall, CurNM is a curriculum learning strategy that dynamically adjusts the learning difficulty of negative samples based on the model’s performance, thereby addressing the prevalent issue of steep difficulty curves in existing methods. Subsequently, we incorporate two components to address the unique characteristics of temporal networks. Firstly, to address positive sparsity, we construct a negative pool for selection, which includes a mix of random samples and historical neighbors, with hard negatives added if the model shows strong learning progress. Starting with a relatively small negative pool also helps maintain computational efficiency. Secondly, we design a negative selection function that leverages temporal-aware disentanglement to capture positive shifts and individual node preferences, which further enhances the model’s granularity.
@@
    Elaborate experiments on 12 temporal networks and 3 TGNNs show that our CurNM can consistently boost the link prediction performance compared to baselines.
@@
    CurNM is highly portable and can be integrated as a plug-in into any TGNN with minimal modifications.
@@ 2.1 Temporal Graph Learning
Existing TGNNs can be classified into discrete-time and continuous-time TGNNs based on their temporal natures [13]. On the one hand, discrete-time TGNNs [8, 22, 25, 27] focus on the sequential dynamics of a series of graph snapshots, where the timespan of a snapshot depends largely on the graph domain. The consequent discrete-time TGNNs include a combination of Recurrent Neural Networks and Graph Neural Networks. DySAT [25], a typical discrete-time TGNN, first learns the structural representations of nodes within each graph snapshot and then aggregates across multiple snapshots using a temporal self-attention layer.
On the other hand, continuous-time TGNNs focus on the sequential interactions of nodes, thereby offering more fine-grained temporal information. Various design paradigms exist for continuous-time TGNNs since temporal networks can be seen as sequential dynamics of either nodes or a graph. JODIE [16] utilizes two mutually updated RNNs to update user and item embeddings, alongside a temporal attention layer that projects user embeddings after a certain time since their last interaction. Conversely, TGN [24] presents a generic framework on continuous temporal networks, simulating both RNNs [16, 30] and TGNNs [33] with its flexible memory mechanism and GNNs. Recent state-of-the-art continuous-time TGNNs [34, 37] have further explored the self-attention mechanism on temporal networks, inspired by the advancement of Transformer networks [2]. However, TGNNs typically employ random and recent negative sampling [24, 33], with little attention given to the specific study of negative sampling strategies within this domain.
@@ 2.2 Hard Negative Sampling
Negative sampling strategies [4, 5, 12, 17, 32, 36, 38, 39] have been abundant in graph representation learning such as collaborative filtering and knowledge graphs. These strategies are traditionally categorized into two types: static and model-aware negative sampling.
Figure 1: An illustration of our CurNM method for temporal networks, which progressively increases the difficulty of sampled negatives. $(a)$ illustrates the pipeline of TGNNs in temporal network link prediction, where our CurNM adds a Negative Pool and a Negative Selection module to the basic random sampling approach. $(b)$ illustrates the construction of a fixed-size, dynamically updated negative pool. $(c)$ shows the selection of negatives from this pool based on the candidates’ disentangled distances and their temporal proximity to the interactions. $(d)$ depicts the final negative samples used for training TGNNs, which include selected negatives from $(c)$ combined with random negatives. $(e)$ illustrates the process of selecting negatives in random sampling, where irrelevant negatives are likely to be chosen at any stage of training.
@@
To the best of our knowledge, ENS [7] is the only negative sampling method specifically designed for temporal networks. ENS selects negative samples by aiming to meet a difficulty target composed of the proportion of historical neighbors and average temporal proximity. While ENS tries to tackle positive sparsity by increasing the difficulty target by a predefined amount each epoch, its heuristic approach fails to adapt to the varying needs of different models and datasets. Moreover, although it addresses positive shift by considering temporal proximity, this method requires evaluating all historical edges to ensure accuracy, thereby consuming substantial resources. Finally, ENS does not account for individual node preferences; it indiscriminately prioritizes all nodes that have historically appeared, without distinguishing historical neighbors, and fails to recognize that interactions between users and items are often driven by certain key factors.
@@ 4 Methodology
Overall, CurNM is a curriculum learning strategy that progressively increases the difficulty of selected negatives based on performance on the validation set, effectively addressing the steep difficulty curves found in existing methods. Within this framework, we further develop two components. Firstly, to tackle positive sparsity, we establish a negative pool consisting of historical neighbors and random negatives, with hard negatives included when the model exhibits sustained improvements. Secondly, we use a temporal-aware disentangling technique to score the information of candidate negatives in the negative pool, emphasizing those that differ from positives only in the most critical dimensions, thus capturing positive shifts and individual preferences. These informative negative samples are then fetched for training TGNNs to ensure their performance and robustness.
@@
3:     Sample a mini-batch $O^{+}_{\text{batch}}$ from $O^{+}$.
@@
6:        Check $\mathcal{C}_{\pi_{e}}$ whether to cache training negatives by Eq. 5.
7:        if not $\mathcal{C}_{\pi_{e}}$ then
@@
4.1 Curriculum Learning
To address the learning difficulty caused by positive sparsity, the curriculum learning strategy adaptively increases the difficulty level of selected negatives based on the learning performance of TGNNs. Additionally, these selected negatives are accompanied by annealing random negatives for robust training of TGNNs.
Adaptive Selected Negatives
Firstly, we introduce a sampling proportion $\pi_{e}$ that self-adjusts based on the model’s performance at $e$-th epoch – shrinking when performance improves and expanding otherwise – thus dynamically controlling the difficulty level of selected negatives. By starting with a large $\pi_{e}$ to include both easy and hard negatives and tightening it only if the model consistently demonstrates improvements, this approach prevents the model from being overwhelmed by overly challenging negatives. Furthermore, we utilize a shrinkage rate $\gamma_{\text{shrink}}$ to control the adjustment steps of $\pi_{e}$. A high $\gamma_{\text{shrink}}$ establishes a steep difficulty curve for model learning, while a low $\gamma_{\text{shrink}}$ sets a flatter curve. This strategy is written as
@@ 4.1 Curriculum Learning
In addition, instead of the traditional method of selecting a set of negative samples for each positive sample, we merge and evaluate all candidate negatives at the minibatch level. As the learning difficulties vary across source nodes, this approach allows us to focus on more informative nodes that present greater difficulty while automatically eliminating nodes that are easier to learn as $\pi_{e}$ shrinks. Technically, we select the top-scored $n_{b}=\pi_{e}\times|b|\times M$ samples for each minibatch $b$ at $e$-th epoch, where $|b|$ denotes the number of source nodes in minibatch $b$, and $M$ represents the size of each source’s negative pool $V^{-}_{u}$.
$\begin{split}O^{-}_{\text{sel}}&=\text{arg top}_{n_{b}}\{\hat{s}_{uv^{-}}|\forall u\in b,v^{-}\in V^{-}_{u}\},\end{split}$ (2)
where $O^{-}_{\text{sel}}$ denotes the set of selected negatives, and $\hat{s}_{uv^{-}}$ is a scoring function between node pairs for negative selection, as defined by Eq. 13.
Annealing Random Negatives
To provide the model with additional space for exploration and adaptation, especially in early epochs, each minibatch also includes $n_{b}$ random negatives $O^{-}_{\text{rand}}$ in addition to $O^{-}_{\text{sel}}$. A weight $\delta_{e}$ is applied to balance these two types. Initially, we assign a higher weight to random negatives to enable the model to understand the basic patterns. As the model gains experience, the emphasis gradually shifts toward the selected, presumably more challenging negatives, enhancing the learning depth. Following this intuition, we define $\delta_{e}=\max\{\pi_{e},\delta_{\text{min}}\}$. $\delta_{\text{min}}$ is a hyper-parameter that controls the minimum influence random negatives should exert, ensuring that both contribute to the learning process. The loss for each minibatch is computed as

### v2

@@ Abstract
Temporal networks are effective in capturing the evolving interactions of networks over time, such as social networks and e-commerce networks. In recent years, researchers have primarily concentrated on developing specific model architectures for Temporal Graph Neural Networks (TGNNs) in order to improve the representation quality of temporal nodes and edges. However, limited attention has been given to the quality of negative samples during the training of TGNNs. When compared with static networks, temporal networks present two specific challenges for negative sampling: positive sparsity and positive shift. Positive sparsity refers to the presence of a single positive sample amidst numerous negative samples at each timestamp, while positive shift relates to the variations in positive samples across different timestamps. To robustly address these challenges in training TGNNs, we introduce Curriculum Negative Mining (CurNM), a model-aware curriculum learning framework that adaptively adjusts the difficulty of negative samples. Within this framework, we first establish a dynamically updated negative pool that balances random, historical, and hard negatives to address the challenges posed by positive sparsity. Secondly, we implement a temporal-aware negative selection module that focuses on learning from the disentangled factors of recently active edges, thus accurately capturing shifting preferences. Finally, the selected negatives are combined with annealing random negatives to support stable training. Extensive experiments on 12 datasets and 3 TGNNs demonstrate that our method outperforms baseline methods by a significant margin. Additionally, thorough ablation studies and parameter sensitivity experiments verify the usefulness and robustness of our approach. Our code is available at https://github.com/zziyue83/CurNM.
@@
[2]organization=Zhejiang Provincial Engineering Research Center for Real-Time SmartTech in Urban Security Governance, School of Computer and Computing Science, Hangzhou City University, city=Hangzhou, state=Zhejiang, country=China
@@ 1 Introduction
Similar to static graph models, TGNNs [30, 39] employ negative sampling techniques to prevent the models from being overwhelmed by a disproportionally high number of negative over positive samples and enhance the models’ efficiency. The most straightforward approaches often adopted are random sampling and recent sampling [30, 39]. However, temporal networks introduce two unique challenges that particularly undermine these approaches: positive sparsity and positive shift. Positive sparsity emerges as the inclusion of a temporal dimension significantly increases the negative-to-positive ratio by the number of timestamps, compared to static graphs. Consequently, it becomes particularly challenging to identify meaningful negative samples for training temporal networks, necessitating either the synthesis of additional positive samples or a more careful selection of negative ones. Positive shift, on the other hand, occurs as the relationships between nodes change over time; for instance, in recommendation systems, users’ preferences are likely to evolve [20]. Ignoring this factor may result in training models on outdated information, particularly in datasets spanning long periods.
@@
To address these challenges, we propose CurNM, a curriculum learning strategy that dynamically adjusts the difficulty of negative samples based on the model’s performance. This approach effectively mitigates the steep learning curves often encountered in existing methods. Within this framework, we introduce two key components tailored to the unique characteristics of temporal networks. First, to handle positive sparsity, we construct a negative pool comprising a mix of random samples and historical neighbors. Hard negatives are introduced as the model demonstrates stronger learning capabilities. Starting with a smaller negative pool also ensures computational efficiency. Second, we develop a negative selection function that incorporates temporal-aware disentanglement to capture positive shifts and individual node preferences, further enhancing the model’s granularity.
@@
    Extensive experiments on 12 temporal network datasets using 3 TGNN models demonstrate that CurNM consistently improves link prediction performance when paired with models that capture fine-grained temporal dynamics. Its effectiveness is especially pronounced on datasets with strong temporal characteristics or significant temporal evolution.
@@
    CurNM introduces minimal computational overhead, is highly portable, and can be integrated as a plug-in into any TGNN with minimal modifications, making it well-suited for practical deployment.
@@ 2.1 Temporal Graph Learning
Existing TGNNs can be classified into discrete-time and continuous-time TGNNs based on their temporal natures [17].
On the one hand, discrete-time TGNNs [6, 25, 28, 31, 40, 43] focus on the sequential dynamics of a series of graph snapshots, where the timespan of a snapshot depends on the specific graph domain. Early approaches like EvolveGCN [28] and DGNN [25] combine recurrent models such as GRU [5] and LSTM [11] with GCNs, integrating temporal memory into graph learning. Transformer-based models like DGT [6] and DySAT [31] extend this by leveraging self-attention to capture both structural dependencies and temporal dynamics in evolving graphs. More recent methods streamline dynamic graph learning using general-purpose strategies. For instance, DyTed [43] uses contrastive learning to disentangle temporal and stable features, and ROLAND [40] introduces a meta-learning framework that transforms static GNNs into dynamic ones.
On the other hand, continuous-time TGNNs [20, 30, 41, 45] focus on the sequential interactions of nodes, thereby offering more fine-grained temporal information. Various design paradigms exist for continuous-time TGNNs since temporal networks can be seen as sequential dynamics of either nodes or a graph. JODIE [20] utilizes two mutually updated RNNs to update user and item embeddings, alongside a temporal attention layer that projects user embeddings after a certain time since their last interaction. Conversely, TGN [30] presents a generic framework on continuous temporal networks, simulating both RNNs [20, 35] and TGNNs [39] with its flexible memory mechanism and GNNs. Recent state-of-the-art continuous-time TGNNs [41, 45] have further explored the self-attention mechanism on temporal networks, inspired by the advancement of Transformer networks [2]. However, TGNNs typically employ random and recent negative sampling [30, 39], with little attention given to the specific study of negative sampling strategies within this domain.
@@ 2.2 Hard Negative Sampling
Negative sampling strategies [4, 8, 15, 21, 38, 44, 46, 48] have been abundant in graph representation learning, such as collaborative filtering and knowledge graphs. These strategies are traditionally categorized into two types: static and model-aware negative sampling.
@@
To the best of our knowledge, ENS [10] is the only negative sampling method specifically designed for temporal networks. ENS selects negative samples by aiming to meet a difficulty target composed of the proportion of historical neighbors and average temporal proximity. While ENS tries to tackle positive sparsity by increasing the difficulty target by a predefined amount each epoch, its heuristic approach fails to adapt to the varying needs of different models and datasets. Moreover, although it addresses positive shift by considering temporal proximity, this method requires evaluating all historical edges to ensure accuracy, thereby consuming substantial resources. Finally, ENS does not account for individual node preferences; it indiscriminately prioritizes all nodes that have historically appeared, without distinguishing historical neighbors, and fails to recognize that interactions are often driven by certain key factors.
2.3 Positive Sampling
Finally, it is worth mentioning another related line of work focused on the selection of positive rather than negative samples. For example, TASER [7] proposes a logit-based strategy that favors positive samples with higher prediction confidence, aiming to reduce noise during training. While TASER may prioritize relevant and timely interactions over outdated ones - thus partially addressing the challenge of positive shift - it offers limited benefit in tackling the issue of positive sparsity in temporal networks. In contrast, CurNM attempts to address both challenges by selecting informative negatives.
@@ 4 Methodology
Overall, CurNM is a curriculum learning strategy that adaptively adjusts the difficulty level of negative samples during training, based on the learning performance of TGNNs. By doing so, it effectively addresses the two unique characteristics of temporal networks and the steep difficulty curves found in existing methods.
To address the issue of positive sparsity, CurNM first constructs a dynamic negative pool composed of historical neighbors and random negatives, with hard negatives added when the model exhibits sustained improvements. From this pool, it then selects an increasingly difficult set of negatives as the model progresses, with the difficulty curve governed by validation performance. The difficulty of candidate negatives is assessed using a temporally-aware disentangling technique that scores them based on their similarity to positive samples. This approach guides the model in learning which factors drive interactions and emphasizes recent node reoccurrence, allowing it to capture positive shifts and individual-level preferences. Finally, these selected negatives are accompanied by annealing random negatives to support robust training of TGNNs.
@@
Figure 1: An illustration of our CurNM method for temporal networks, which progressively increases the difficulty of sampled negatives. $(a)$ shows the pipeline of a single-layer TGNN, where CurNM augments basic random sampling (weighted by $\delta_{e}$) with selected hard negatives (weighted by $1-\delta_{e}$) using a Negative Pool and a Negative Selection module. $(b)$ illustrates the construction of a fixed-size, dynamically updated negative pool. $(c)$ demonstrates the selection of negatives from this pool based on disentangled factor distances and their temporal proximity to the corresponding positive sample. $(d)$ presents the annealing random negatives, matched in size to the selected negatives in $(c)$, which are used together during training to ensure a smoother learning curve.
@@
3:     Sample a mini-batch $O^{+}_{b}$ from $O^{+}$.
@@
6:        Check $\mathcal{C}_{e}$ whether to cache training negatives by Eq. 2.
7:        if not $\mathcal{C}_{e}$ then
@@
4.1 Dynamic Negative Pool
First, we maintain a negative pool that dynamically balances random, historical, and hard negatives according to the model’s learning performance.
To ensure efficiency, we start the entire process by sampling a negative pool $V^{-}_{u}$ of fixed size $M$ for each positive sample $(u,v^{+},t)$. During the initial epochs, we sample $V^{-}_{u}$ as a mixture of historical neighbors and random negatives, a simple approach to prevent overcrowding with easy negatives. This process can be written as
$V^{-}_{u}=V^{\text{hist}}_{u}\cup V^{\text{rand}}_{u},$ (1)
We increase the difficulty level once the model has learned the fundamentals. To track this learning progress, we define a self-adjusting proportion $\pi_{e}$ based on the model’s performance at epoch $e$, which decreases as performance improves and increases otherwise. [^1] Notably, $\pi_{e}$ also governs the sampling proportion of negatives; its design is detailed in Section 4.2. When $\pi_{e}$ drops below a predefined threshold $\tau$, the advanced stage of the negative pool is activated. This mechanism ensures that more challenging negatives are introduced only when the model is sufficiently prepared. The indicator $\mathcal{C}_{e}$ for this advanced stage is determined as follows:
$\mathcal{C}_{e}=\mathbf{1}(\pi_{e}\leq\tau).$ (2)
In this advanced stage, we implement a hard negative cache $V^{\text{hard}}_{u}$ to store a subset of $V^{-}_{u}$ from the current epoch, prioritizing likely false positives and samples that exhibit limited learning progress. Technically, these are negatives with consistently high probability scores and low variance across epochs. Thus, cached negatives are sampled as
$V^{\text{hard}}_{u}\subset V^{-}_{u},\quad\text{sampled proportionally to}\quad p_{uv}-\alpha_{e}\cdot\mathrm{std}(P_{uv}),$ (3)
where $p_{uv}$ denotes the probability score of node pair $(u,v)$ at the current epoch, $P_{uv}$ records the historical scores within the last five epochs, and std($\cdot$) is set to 00 when results are not available. $\alpha_{e}$ is a hyper-parameter that controls the importance of prediction standard deviation at the $e$-th epoch. To manage potential instability in early learning stages, we initially assign a modest weight to the standard deviation and gradually increase it. Specifically, we use $\alpha_{e}=\alpha_{\text{max}}\times\min(\frac{e}{E^{\prime}},1)$, where $E^{\prime}$ is the epoch after which the weight remains constant, and $\alpha_{\text{max}}$ is the maximum weight that standard deviations could contribute to the performance metrics.
We set $|V^{\text{hard}}_{u}|=\frac{M}{2}$, meaning that the hard negative cache accounts for half of the negative pool in the next epoch. The remaining half continues to be composed of historical neighbors and random samples, maintaining the original sampling proportion. This split ensures that the model is equally exposed to both poorly learned negatives and previously unseen ones. Therefore, the new updating strategy for each negative pool $V^{-}_{u}$ is written as
$V^{-}_{u}=V^{\text{hard}}_{u}\cup V^{\text{hist}}_{u}\cup V^{\text{rand}}_{u},$ (4)
4.2 Adaptive Negative Selection
Given the negative pool, we select the top $\pi_{e}$ fraction of negative samples for model training. Section 4.2.1 explains how the sampling proportion $\pi_{e}$ is determined, and Section 4.2.2 describes the method used to compute the scores $s_{uv^{-}}$, which are used to rank the negative samples.
4.2.1 Sampling Proportion
As mentioned above, the sampling proportion $\pi_{e}$ is dynamically adjusted based on the model’s performance. Specifically, to prevent the model from being overwhelmed by overly challenging examples, we start with a relatively large $\pi_{e}$, allowing the inclusion of both easy and hard negatives. As the model improves, this sampling proportion is gradually reduced, exposing the model to increasingly harder negatives. Conversely, if performance plateaus or degrades, $\pi_{e}$ can be relaxed to include easier examples again. Furthermore, we utilize a shrinkage factor $\gamma_{\text{shrink}}$ to control the adjustment steps of $\pi_{e}$. A high $\gamma_{\text{shrink}}$ establishes a steep difficulty curve for model learning, while a low $\gamma_{\text{shrink}}$ sets a flatter curve. This strategy is written as
@@ 4.2.1 Sampling Proportion
In addition, instead of the traditional method of selecting a set of negative samples for each positive sample, we merge and evaluate all candidate negatives at the minibatch level. As the learning difficulties vary across source nodes, this approach allows us to focus on more informative nodes that present greater difficulty while automatically eliminating nodes that are easier to learn as $\pi_{e}$ shrinks. Collectively, this strategy can be formulated as:
$\begin{split}O^{-}_{\text{sel}}=\text{arg\,top}^{N_{b}}_{\left\{(u,v^{-},t)\mid(u,v^{\prime},t)\in O^{+}_{b},v^{-}\in V^{-}_{u}\right\}}s_{uv^{-}},\end{split}$ (6)
where $O^{-}_{\text{sel}}$ denotes the set of selected negatives, and $s_{uv^{-}}$ is the scoring function used to rank the negative samples, as defined in Eq. 12. The operator $\text{arg\,top}^{N_{b}}_{\left\{(u,v^{-},t)\mid(u,v^{\prime},t)\in O^{+}_{b},v^{-}\in V^{-}_{u}\right\}}$ selects the top $N_{b}$ pairs $(u,v^{-},t)$ with the highest scores. Here, $N_{b}=\pi_{e}\times|b|\times M$, where $|b|$ denotes the number of edges in the minibatch $b$ and $M$ is the number of candidate negatives per edge (i.e., the size of the negative pool $V^{-}_{u}$).
4.2.2 Negative Ranking
We rank negative samples from the negative pool $V^{-}_{u}$ based on their proximity to the corresponding positive samples. To compute this proximity, we disentangle irrelevant factors and enhance relevant factors using time embeddings.
Irrelevant Factor Disentanglement with Temporal Enhancement
Recognizing that temporal networks often exhibit structured and evolving interaction patterns and thus not all information is equally valuable, we aim to decouple the relevant factors driving interactions from the irrelevant ones. To achieve this, we adopt a disentanglement method inspired by DENS [21] and further integrate temporal modeling to better capture the dynamic behaviors inherent in temporal networks.
Following DENS [21], we first construct a positive gate for each positive sample $(u,v^{+},t)$. This gate applies a linear transformation to the concatenated embeddings of the source node $h_{u}$ and the destination node $h_{v^{+}}$ to isolate the relevant factor representation $h_{v^{+}}^{r}$ of the destination node $v^{+}$:
$h_{v^{+}}^{r}=h_{v^{+}}\odot\sigma(W_{p}[h_{u}\,||\,h_{v^{+}}]+b_{p}),$ (7)
where $W_{p}$ and $b_{p}$ are trainable parameters shared across all positive samples, $\sigma$ denotes the sigmoid function, and $\odot$ is the element-wise product.
To better capture the unique temporal characteristics of temporal networks, we further refine the ranking process to prioritize negative samples that are either historically relevant or recently active - samples we refer to as having higher temporal proximity. These negatives often exhibit features previously favored by the source node, making them more difficult to distinguish from true positives.
To achieve this, we incorporate Time2Vec [16], a time embedding method designed to capture complex temporal patterns such as recurrent and periodic behaviors. Specifically, the Time2Vec encoding we use for each time point $\tau$, denoted as $\mathbf{t2v}(\tau)$, is defined as:
$\mathbf{t2v}(\tau)[i]=\begin{cases}W^{t2v}_{i}\tau+b^{t2v}_{i},&\text{if }i=0,\\ \sin(W^{t2v}_{i}\tau+b^{t2v}_{i}),&\text{if }1\leq i\leq k,\end{cases}$
where $\mathbf{t2v}(\tau)[i]$ denotes the $i$-th element of $\mathbf{t2v}(\tau)$, and $W^{t2v}_{i}$ and $b^{t2v}_{i}$ are learnable parameters. The dimension $k$ is chosen to match the dimensionality of the node embeddings.
We first use Time2Vec to encode the timestamp of each positive sample. The resulting time-enhanced factor embedding is computed as:
$\hat{h}_{v^{+}}^{r}=h_{v^{+}}^{r}\odot z(\mathbf{t2v}(t_{uv^{+}})),$ (8)
where $t_{uv^{+}}$ is the timestamp of the positive sample $(u,v^{+},t_{uv^{+}})$, and $z(\cdot)$ denotes a normalization function that prevents the temporal signal from dominating the node representation.
This factor representation $\hat{h}_{v^{+}}^{r}$, along with the embedding $h_{v^{-}}$ of a candidate negative $v^{-}\in V_{u}^{-}$, is then used to isolate the relevant factor $h_{v^{-}}^{r}$ of the negative:
$h_{v^{-}}^{r}=h_{v^{-}}\odot\sigma(W_{n}[\hat{h}_{v^{+}}^{r}\,||\,h_{v^{-}}]+b_{n}),$ (9)
where $W_{n}$ and $b_{n}$ are trainable parameters shared across all candidate negatives.
Finally, we use Time2Vec to encode two timestamps that reflect temporal proximity: $t_{uv^{-}}$, the timestamp of the historical interaction between $u$ and $v^{-}$, and $t_{v^{-}}$, the timestamp of the most recent occurrence of $v^{-}$ before the current interaction. If the negative sample is not a historical neighbor or has never appeared before, default values of 0 are passed into Time2Vec. The final factor embedding of the negative sample is then computed as:
$\hat{h}_{v^{-}}^{r}=h_{v^{-}}^{r}\odot z(\mathbf{t2v}(t_{uv^{-}})\oplus\mathbf{t2v}(t_{v^{-}})),$ (10)
where $\oplus$ denotes element-wise addition.
Factors other than relevant factors are denoted as irrelevant factors:
$\hat{h}_{v^{+}}^{ir}=h_{v^{+}}\ominus\hat{h}_{v^{+}}^{r},\quad\hat{h}_{v^{-}}^{ir}=h_{v^{-}}\ominus\hat{h}_{v^{-}}^{r},$ (11)
where $\hat{h}_{v^{+}}^{ir}$ and $\hat{h}_{v^{-}}^{ir}$ denote irrelevant factors of the destination node $v^{+}$ and the candidate negative $v^{-}$, respectively, and $\ominus$ is element-wise subtraction.
Scoring Function
As our goal is to refine the decision boundary as the model progresses, we aim to train it with negatives that increasingly resemble positive samples. We achieve this in two phases. In the first phase, the model focuses on learning which factors are important and how they influence interactions. This involves initially aligning irrelevant factors, encouraging the model to filter out noise and identify signals. In the second phase, we increase the difficulty by also aligning the relevant factors, pushing the model to make finer-grained distinctions between positive and negative samples. Technically, we define the overall scoring function used to select $O^{-}_{\text{sel}}$ as:
$s_{uv^{-}}=-\beta_{e}\times|s_{r,r}^{+,-}|-(2-\beta_{e})\times|s_{ir,ir}^{+,-}|,$ (12)
where $s_{g_{1},g_{2}}^{\text{sgn}_{1},\text{sgn}_{2}}=\text{MLP}(h_{u},\hat{h}_{v^{\text{sgn}_{1}}}^{g_{1}})-\text{MLP}(h_{u},\hat{h}_{v^{\text{sgn}_{2}}}^{g_{2}})$ represents the relative score between two factor-specific representations. $\text{MLP}(\cdot,\cdot)$ computes a probability score from the given embeddings. [^2] Whether the two inputs are added or concatenated depends on the architecture of the underlying TGNN. A common implementation uses $\texttt{ReLU}(h_{u}+h_{v})$ followed by a linear projection. We refer to $|s_{r,r}^{+,-}|$ as the relevant distance, as it measures the score difference between the relevant factors of the positive and negative samples. Similarly, $|s_{ir,ir}^{+,-}|$ is termed the irrelevant distance. The hyperparameter $\beta_{e}$ controls the balance between relevant and irrelevant factors at epoch $e$. We set $\beta_{e}=\min\left(\frac{e}{E^{\prime\prime}},1\right)$, where $E^{\prime\prime}$ is the epoch after which both factors contribute equally.
Supervised Disentanglement Learning
where the subtraction is added to perform maximization.
4.3 Annealing Random Negatives
To provide the model with additional space for exploration and adaptation, especially in early epochs, each minibatch also includes $N_{b}$ random negatives, forming the set $O^{-}_{\text{rand}}$, in addition to the selected negative set $O^{-}_{\text{sel}}$.
A weight $\delta_{e}$ is applied to balance these two types. Initially, we assign greater weight to randomly sampled negatives, allowing the model to acquire foundational knowledge through standard sampling mechanisms. As training progresses, this knowledge is then leveraged to handle selected, presumably more challenging negatives. This process is analogous to a semi-supervised learning approach [22, 37, 47], where the model first learns from labeled data and later uses that knowledge to benefit from unlabeled data.
To implement this annealing behavior, we define $\delta_{e}=\max\{\pi_{e},\delta_{\text{min}}\}$. By linking $\delta_{e}$ to $\pi_{e}$, we align the influence of random negatives with the model’s performance: when the model’s performance improves, $\pi_{e}$ decreases, and consequently, $\delta_{e}$ also decreases, gradually reducing the role of random negatives. $\delta_{\text{min}}$ is a hyper-parameter that controls the minimum influence random negatives should exert, ensuring both contribute to the learning process. The loss for each minibatch is computed as